OpenStudy (anonymous):

The energy (and cost) needed to operate an air conditioner is proportional to the volume of the space that is being cooled. It costs ZAP Electronics about $125.00 per day to run an air conditioner in its small rectangular warehouse. The company's large warehouse, a few blocks away, is 2.5 times as long, wide, and high as the small warehouse. Estimate the daily cost of cooling the large warehouse with the same model of air conditioner. Round your answer to the nearest dollar.

OpenStudy (anonymous):

Just show me step by step please.

OpenStudy (anonymous):

Well, just as a quick'n'dirty, I'd first imagine the little building as being a 1 x 1 x 1 = 1^3 = 1cu box. That would make the larger building 2.5 x 2.5 x 2.5 = 2.5^3 = 15.625cu So any building that's 2.5 times larger than another building has a volume that's 15.625 times as large as the smaller one.

OpenStudy (anonymous):

Makes sense. So I would then multiply that number by the money to get how much money it would cost to do the 2.5 box?

OpenStudy (anonymous):

This board is driving me CRAZY with it's timeouts, dropped connections and it's "guarding new features" crapola! Anyhow, yes... :-) Since the problem states that the cost is proportional, I would simply multiply the cost by the same ratio as the volume.

OpenStudy (anonymous):

So, Yes, what I am doing above is correct, and that's the answer? Great! Thanks for your help! :D

OpenStudy (anonymous):

No problem!! Glad I could help, Claw! And thank you! :-)